OpenStudy (anonymous):

13.Find the geometric mean of 3 and 7. √21 2√6 2√7 21 14.Find the geometric mean of 18 and 8. 22 12 17 32

OpenStudy (whpalmer4):

The geometric mean between two numbers a and b is the number x such that \(a:x\) is as \(x:b\). Written as fractions, \[\frac{a}{x} = \frac{x}{b}\] To find x, cross-multiply to get \[x^2 = ab\]and take the square root of both sides.